The lengths of the perpendiculars from the focus and the extremities of a focal chord of a parabola on the tangent at the vertex form
A. HP
B. GP
C. AP
 
Answer is GP

Grade:11

1 Answers

Aditya Gupta
6 years ago
consider a parabola y^2=4ax; a>0, with focus (a,0) and tangent to the vertex being x=0
now, let a focal chord intersect the parabola at (at^2, 2at) and (ap^2, 2ap)
as this chord passes thru focus, we have
[2ap-2at]/[ap^2-at^2]=2ap/[ap^2-a]
or pt= –1
or p^2t^2=1
now, the lengths from the tangent at vertex are:
at^2, a, ap^2
we see that a/at^2=ap^2/a
so, the lengths are in GP